# Who Gets Cited Most? Benchmarking Long-Context Language Models on Scientific Articles

### v1

@@ Appendix F Experimental Details
Appendix G Additional Results
| Full-text Articles | Database TablesΔ | ||||||||
| Model | CLen | 64K | 128K | 512K | 1M | 64K | 128K | 512K | 1M |
@@ Appendix G Additional Results
Appendix H Reasoning Traces for Reinforcement Learning
The prompt used by GRPO to generate reasoning traces is given in Figure 7.
Appendix I Details for Model Failure Analyses
I.1 Analyses on Model Answers
To gain deeper insights into model behavior, we analyzed the failure patterns exhibited in model-generated answers on SciTrek. We randomly selected 60 instances in the input length of 128K that were hard for all zero-shot and supervised models (10 instances per skill in Table 4). We find that models tend to output “NULL” particularly on samples that require the skill of filtering. As shown in Table 14, most open-weight models tend to randomly output “NULL” for any sample. This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context.141414Our testing instruction prompted models to generate “NULL” when they could not get the answer from the context. In addition, we find that models fail to follow the specified format for the answers, especially for sorting-related questions, as shown in Table 15. For example, models tend to generate author lists when sorted author numbers are requested and output lists when aggregates are requested. Interestingly, GPT-4.1 produces incorrectly formatted answers far more frequently than all other models. We further observe that some models tend to produce partial answers when lists are requested, especially for tasks that require aggregation. As shown in Table 16, our GRPO-based model tends to generate partial answers more than other models.
@@ I.1 Analyses on Model Answers
| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
| Model | +Aggregation | +Sorting | Filtering | |||
@@
| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
| Model | +Aggregation | +Sorting | Filtering | |||
@@
| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
| Model | +Aggregation | +Sorting | Filtering | |||
@@
| Filtering | Filtering | Filtering | Relational | Average | |
| Model | +Aggregation | +Sorting | Filtering | ||
@@
I.2 Example Reasoning Traces from GRPO-based Model
We randomly sampled reasoning traces for three SQL query templates from the GRPO-based model. The reasoning traces for different sizes of contexts are presented in Table LABEL:tab:example_reasoning_traces.
[…]
Question
Context 64K |
What is the smallest number of authors for articles that have more than 13 references? |
[…]
To solve this problem, we need to follow these steps:
1. Identify articles with more than 13 references.
2. Determine the number of authors in each of those articles.
3. Find the smallest number of authors among these articles.
Let’s go through each step in detail:
### Step 1: Identify articles with more than 13 references
From the provided articles, we need to count the references in each article:
1. **”Best Constants in Young’s Inequality, Its Converse, and Its Generalization to More than Three Functions”**
- References: 13 (listed in the reference section)
2. **”Optical detection of strain and doping inhomogenieties in single layer MoS2”**
- References: 62 (listed in the reference section)
3. **”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”**
- References: 65 (listed in the reference section)
4. **”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels”**
- References: 32 (listed in the reference section)
Only the articles titled ”Optical detection of strain and doping inhomogenieties in single layer MoS2”, ”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”, and ”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels” have more than 13 references.
### Step 2: Determine the number of authors in each of these articles
Now, let’s count the authors in each of these articles:
1. **”Optical detection of strain and doping inhomogenieties in single layer MoS2”**
- Authors: A. Michail, N. Delikoukos, J. Parthenios, C. Galiotis, K. Papagelis (5 authors)
2. **”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”**
- Authors: Akiko Ueda, Yijin Zhang, Nobuyuki Sano, Hiroshi Imamura, Yoshihiro Iwasa (5 authors)
3. **”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels”**
- Authors: Will Yeadon, Tom Hardy (2 authors)
### Step 3: Find the smallest number of authors
Among the articles with more than 13 references, the smallest number of authors is 2.
[…]
Question
Context 128K |
What is the smallest number of authors for articles that have more than 25 references? |
[…]
To solve this problem, we need to follow the steps outlined in the question. Let’s go through each step in detail.
### Step 1: Identify Articles with More Than 25 References
First, we need to go through each provided article and determine which ones have more than 25 references. We’ll count the words in the titles using spaces as delimiters to ensure accuracy.
#### Article Titles and Reference Counts:
1. **Article Title:** ”A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE”
- Words: Count spaces to determine words.
A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE
Spaces: 8 words (A, SYSTEMATIC, EVALUATION, OF, LARGE, LANGUAGE, MODELS, OF, CODE)
- Number of references: 27 (as given in the problem statement).
This article has more than 25 references.
2. **Article Title:** ”Reputation-based partner choice promotes cooperation in social networks”
- Words: Count spaces to determine words.
Reputation-based partner choice promotes cooperation in social networks
Spaces: 9 words (Reputation-based, partner, choice, promotes, cooperation, in, social, networks).
- Number of references: 26 (as given in the problem statement).
This article has more than 25 references.
<...we cut it down to save space...>
### Step 2: Determine the Smallest Number of Authors in Articles with More Than 25 References
Now that we’ve identified the articles with more than 25 references, let’s count the authors in each of these articles.
1. **Article Title:** ”A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE”
- Authors: Frank F. Xu, Uri Alon, Graham Neubig, Vincent J. Hellendoorn (4 authors).
2. **Article Title:** ”Reputation-based partner choice promotes cooperation in social networks”
- Authors: Feng Fu, Christoph Hauert, Martin A. Nowak, Long Wang (4 authors).
3. **Article Title:** ”Efficient Multimodal Sampling via Tempered Distribution Flow”
- Authors: Yixuan Qiu, Xiao Wang (2 authors).
4. **Article Title:** ”Ultralow contact resistance between semimetal and monolayer semiconductors”
- Authors: Pin-Chun Shen, Cong Su, Yuxuan Lin, Ang-Sheng Chou, Chao-Ching Cheng, Ji-Hoon Park, Ming-Hui Chiu, Ang-Yu Lu, Hao-Ling Tang, Mohammad Mahdi Tavakoli, Gregory Pitner, Xiang Ji, Zhengyang Cai, Nannan Mao, Jiangtao Wang, Vincent Tung, Ju Li, Jeffrey Bokor, Alex Zettl, Chih-I Wu, Tomás Palacios, Lain-Jong Li, Jing Kong (11 authors).
5. **Article Title:** ”Normal forms for singularities of pedal curves produced by non-singular dual curve germs in Sn”
- Authors: Takashi Nishimura (1 author).
### Step 3: Identify the Smallest Number of Authors
[…]
### Final Answer:
[…]
Question
Context 512K |
What is the smallest number of authors for articles that cite more than 84 references? |
[…]
To solve this problem, we need to follow these steps:
1. Identify articles that cite more than 84 references.
2. Determine the number of authors in each qualifying article.
3. Identify the smallest number of authors among these articles.
4. Provide the final answer in the required format.
Let’s go through each step in detail.
### Step 1: Identify articles that cite more than 84 references
First, we will examine each provided article to determine which ones have more than 84 references.
1. **Article Title:** ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface”
Number of words in title: Counting spaces, we have ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface” → Words are: Macroscopic, model, and, statistical, model, to, characterize, electromagnetic, information, of, a, digital, coding, metasurface. (12 words).
Number of references: 85 (as stated in the problem).
2. **Article Title:** ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set”
Number of words in title: ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set” → Words are: Worst-case, VaR, and, robust, portfolio, optimization, with, interval, random, uncertainty, set. (11 words).
Number of references: 84 (as stated in the problem).
<...we cut it down to save space...>
Now we have identified the following articles that cite more than 84 references:
- ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface” (85 references).
- ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set” (84 references).
- ”Direct Transmission of Digital Message via Programmable Coding Metasurface” (82 references).
- ”Bayesian experimental design for linear elasticity” (84 references).
- ”Metaverse for Social Good: A University Campus Prototype” (84 references).
- ”Alloying effect-induced electron polarization drives nitrate electroreduction to ammonia” (84 references).
### Step 2: Determine the number of authors in each qualifying article
Next, we will count the authors in each of these identified articles.
1. **”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface”**
Authors listed: Rui Wen Shao, Wensheng Gan (listed as corresponding author), Philip S. Yu (listed as corresponding author).
Number of authors: 3.
2. **”Worst-case VaR and robust portfolio optimization with interval random uncertainty set”**
Authors listed: Wei Chen, Shaohua Tan, Dongqing Yang.
Number of authors: 3.
3. **”Direct Transmission of Digital Message via Programmable Coding Metasurface”**
Authors listed: Tie Jun Cui, Shuo Liu, Guo Dong Bai, Qian Ma.
Number of authors: 4.
4. **”Bayesian experimental design for linear elasticity”**
Authors listed: Haihan Duan, Jiaye Li, Sizheng Fan, Zhonghao Lin, Xiao Wu, Wei Cai (listed as corresponding author).
Number of authors: 6.
<...we cut it down to save space...>
### Step 3: Identify the smallest number of authors among these articles
Among the identified articles that cite more than 84 references, the smallest number of authors is 3.
### Final Answer
[…]
Question
Context 1M |
What is the smallest number of authors for articles that cite more than 51 references? |
[…]
Question
Context 64K |
How many different articles has Shi Jin been listed as the first author of? |
[…]
To determine how many different articles Shi Jin has been listed as the first author of, we need to go through each article provided in the documents and check his position as an author. Let’s examine each article one by:
1. **Broadband diffusion of terahertz waves by multi-bit coding metasurfaces**:
- Authors: Li-Hua Gao, Qiang Cheng, Jing Yang, Shao-Jie Ma, Jie Zhao, Shuo Liu, Hai-Bing Chen, Qiong He, Wei-Xiang Jiang, Hui-Feng Ma, Qi-Ye Wen, Lan-Ju Liang, Biao-Bing Jin, Wei-Wei Liu, Lei Zhou, Jian-Quan Yao, Pei-Heng Wu, Tie-Jun Cui
- Shi Jin is not listed as the first author here.
2. **Large Intelligent Surface-Assisted Wireless Communication Exploiting Statistical CSI**:
- Authors: Yu Han, Wankai Tang, Shi Jin, Chao-Kai Wen, Xiaoli Ma
- Shi Jin is listed as the fourth author but not the first.
3. **Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design**:
- Authors: Qingqing Wu, Rui Zhang
- Shi Jin is not listed as an author here.
4. **Wireless Communication Based on Information Metasurfaces**:
- Authors: Jun Yan Dai, Wankai Tang, Ming Zheng Chen, Chi Hou Chan, Qiang Cheng, Shi Jin, Tie Jun Cui
- Shi Jin is listed as the seventh author but not the first.
After examining all provided articles, Shi Jin has not been listed as the first author in any of them.
Thus, the number of different articles Shi Jin has been listed as the first author of is .
[…]
Question
Context 128K |
How many distinct articles has Tang been the first author of? |
[…]
To solve this problem, we need to follow a systematic approach by analyzing each article provided in the documents. Let’s go through each step in detail.
### Step 1: Identify Articles Written by Tang as the First Author
First, we need to examine each article to determine if Tang is listed as the first author. We will go through each article one by:
#### Article 1: ”The humping phenomenon during high speed gas metal arc welding”
Authors listed: T. C. Nguyen, D. C. Weckman, D. A. Johnson, H. W. Kerr (D. C. Weckman is the corresponding author).
Tang is not listed as an author in this article.
#### Article 2: ”Physics of Humping Formation in Laser Powder Bed Fusion”
Authors listed: C. Tang, K. Q. Le, C. H. Wong (C. Tang is the corresponding author).
Tang is listed as the first author in this article.
#### Article 3: ”Keyhole fluctuation and pore formation mechanisms during laser powder bed fusion additive manufacturing”
Authors listed: Yuze Huang 1,2, Tristan G. Fleming 3, Samuel J. Clark 1,2,4, Sebastian Marussi 1,2, Kamel Fezzaa 4, Jeyan Thiyagalingam 5, Chu Lun Alex Leung 1,2, Peter D. Lee 1,2 (Yuze Huang is the first author).
Tang is not listed as an author in this article.
#### Article 4: ”Unveiling mechanisms and onset threshold of humping in high-speed laser welding”
Authors listed: Zen-Hao Lai 1, Siguang Xu 2, Samuel J. Clark 3, Kamel Fezzaa 3, Jingjing Li 1,4 (Zen-Hao Lai is the first author).
Tang is not listed as an author in this article.
#### Article 5: ”Understanding of humping formation and suppression mechanisms using the numerical simulation”
Authors listed: Dongsheng Wu a,b,c, Xueming Hua a,b,∗, Dingjian Ye a,b, Fang Li a,b (Dongsheng Wu is the first author).
Tang is not listed as an author in this article.
#### Article 6: ”Investigation of the humping formation in the high power and high speed laser welding”
Authors listed: Yuewei Ai a,b, Ping Jia a,b,∗, Chunming Wang b, Gaoyang Mi b, Shaoning Geng a, Wei Liu a, Chu Han a (Yuewei Ai is the first author).
Tang is not listed as an author in this article.
### Step 2: Count Distinct Articles Written by Tang as the First Author
From the analysis above, only one article lists Tang as the first author:
- ”Physics of Humping Formation in Laser Powder Bed Fusion” by C. Tang, K. Q. Le, C. H. Wong.
Thus, Tang has been the first author in exactly one distinct article.
### Final Answer
[…]
Question
Context 512K |
How many different articles was Connor M. Brennan the first author of? |
[…]
To solve this problem, we need to follow the steps outlined in the question. Let’s go through each step carefully and determine the final answer.
### Step 1: Identify Relevant Articles
First, we need to scan through the provided articles to identify which ones have Connor M. Brennan as the first author. This will help us determine how many different articles he authored.
#### Article Titles and Authors:
1. **”The asymmetric effects of monetary policy on stock price bubbles”** by Christophe Blot, Paul Hubert, Fabien Labondance.
2. **”Deconstructing monetary policy surprises: the role of information shocks”** by Marek Jarociński, Peter Karadi.
3. **”Nominal Rigidities and the Dynamic Effects of a Shock to Monetary Policy”** by Lawrence J. Christiano, Martin Eichenbaum, Charles Evans.
4. **”Risk Appetite and the Risk-Taking Channel of Monetary Policy”** by Michael D. Bauer, Ben S. Bernanke, Eric Milstein.
5. **”The Macroeconomic Effects of the Federal Reserve’s Conventional and Unconventional Monetary Policies”** by Eric T. Swanson.
6. **”Federal Reserve Information and the Behavior of Interest Rates”** by Christina D. Romer, David H. Romer.
7. **”A Reassessment of Monetary Policy Surprises and High-Frequency Identification”** by Michael D. Bauer, Eric T. Swanson.
8. **”An Alternative Explanation for the ’Fed Information Effect’”** by Michael D. Bauer, Eric T. Swanson.
9. **”Monetary Policy Shocks: Data or Methods?”** by Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
10. **”High Asymmetric Effects of Monetary Policy on Stock Price Bubbles”** by Christophe Blot, Paul Hubert, Fabien Labondance (Note: This is a duplicate title and should be ignored).
From this list, only one article has Connor M. Brennan as the first author:
- **”Monetary Policy Shocks: Data or Methods?”** by Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
### Step 2: Verify Connor M. Brennan as the First Author
Now, let’s confirm that Connor M. Brennan is indeed the first author of the identified article:
- The article title is ”Monetary Policy Shocks: Data or Methods?”
- The authors listed are Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
- Connor M. Brennan is listed first, confirming he is the first author.
### Step 3: Count Words in the Title Using Spaces
Next, we need to count the words in the title ”Monetary Policy Shocks: Data or Methods?” using spaces as delimiters. Let’s list each word:
1. Monetary
2. Policy
3. Shocks:
4. Data
5. or
6. Methods?
There are 6 words in the title.
### Final Answer
[…]
Question
Context 1M |
How many different articles has Saurabh Bodhe authored as the first author? |
[…]
Question
Context 64K |
How many different articles has Feng Miao authored? |
[…]
To determine how many different articles Feng Miao has authored, we need to go through each article provided and identify Feng Miao’s involvement in each one. Let’s analyze each article one by step.
### Article 1: ”Hopping transport through defect-induced localized states in molybdenum disulphide”
Authors listed: Hao Qiu, Tao Xu, Zilu Wang, Wei Ren, Haiyan Nan, Zhenhua Ni, Qian Chen, Shijun Yuan, Feng Miao, Fengqi Song, Gen Long, Yi Shi, Litao Sun, Jinlan Wang, Xinran Wang.
Feng Miao is listed as an author in this article.
### Article 2: ”Approaching the quantum limit in two-dimensional semiconductor contacts”
Authors listed: Weisheng Li, Xiaoshu Gong, Zhihao Yu, Liang Ma, Wenjie Sun, Si Gao, Çağıl Köroğlu, Wenfeng Wang, Lei Liu, Taotao Li, Hongkai Ning, Dongxu Fan, Yifei Xu, Xuecou Tu, Tao Xu, Litao Sun, Wenhui Wang, Junpeng Lu, Zhenhua Ni, Jia Li, Xidong Duan, Peng Wang, Yuefeng Nie, Hao Qiu, Yi Shi, Eric Pop, Jinlan Wang, Xinran Wang.
Feng Miao is not listed as an author in this article.
### Article 3: ”Origins of genuine Ohmic van der Waals contact between indium and MoS2”
Authors listed: Bum-Kyu Kim, Tae-Hyung Kim, Dong-Hwan Choi, Hanul Kim, Kenji Watanabe, Takashi Taniguchi, Heesuk Rho, Ju-Jin Kim, Yong-Hoon Kim, Myung-Ho Bae.
Feng Miao is not listed as an author in this article.
### Article 4: ”Van der Waals contacts between three-dimensional metals and two-dimensional semiconductors”
Authors listed: Yan Wang, Jong Chan Kim, Ryan J. Wu, Jenny Martinez, Xiuju Song, Jieun Yang, Fang Zhao, Andre Mkhoyan, Hu Young Jeong, Manish Chhowalla.
Feng Miao is not listed as an author in this article.
### Conclusion
Based on the analysis, Feng Miao has authored only one article, which is ”Hopping transport through defect-induced localized states in molybdenum disulphide.”
[…]
Question
Context 128K |
How many different articles has JIAYI SUN authored? |
[…]
To solve this problem, we need to follow these steps:
1. Identify articles authored by JIAYI SUN.
2. Count the number of distinct articles authored by JIAYI SUN.
3. Provide the final answer in the required format.
Let’s go through each step in detail:
### Step 1: Identify articles authored by JIAYI SUN
From the provided articles, we need to examine each one to determine if JIAYI SUN is listed as an author. Let’s review each article:
1. **Article Title:** ”Solipsis: A Decentralized Architecture for Virtual Environments”
Authors listed: Davide Frey, Jérôme Royan, Romain Piegay, Anne-Marie Kermarrec, Emmanuelle Anceaume, Fabrice Le Fessant, and Jiayi Sun.
JIAYI SUN is present here.
2. **Article Title:** ”Metaverse: Survey, Applications, Security, and Opportunities”
Authors listed: Jiayi Sun, Wensheng Gan, Han-Chieh Chao, and Philip S. Yu.
JIAYI SUN is present here.
3. **Article Title:** ”Big Data Meets Metaverse: A Survey”
Authors listed: Jiayi Sun, Wensheng Gan, Zefeng Chen, Junhui Li, and Philip S. Yu.
JIAYI SUN is present here.
4. **Article Title:** ”A Dynamic Resource Allocation Framework for Synchronizing Metaverse with IoT Service and Data”
Authors listed: Yue Han, Dusit Niyato, Cyril Leung, Chunyan Miao, and Jiayi Sun.
JIAYI SUN is present here.
Thus, JIAYI SUN has authored four distinct articles.
### Step 2: Count the number of distinct articles authored by JIAYI SUN
From the analysis above, we have identified the following articles authored by JIAYI SUN:
1. ”Solipsis: A Decentralized Architecture for Virtual Environments”
2. ”Metaverse: Survey, Applications, Security, and Opportunities”
3. ”Big Data Meets Metaverse: A Survey”
4. ”A Dynamic Resource Allocation Framework for Synchronizing Metaverse with IoT Service and Data”
[…]
### Final Answer
[…]
Question
Context 512K |
How many different articles has Qiang Cheng authored? |
[…]
To solve this problem, we need to follow a systematic approach by analyzing each provided article to identify instances where Qiang Cheng is listed as an author. Let’s go through each step in detail.
### Step 1: Identify Articles Mentions Qiang Cheng as an Author
First, we will examine each article title and author list to determine which articles include Qiang Cheng as an author.
1. **Article Title:** ”Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design”
[…]
2. **Article Title:** ”Broadband continuous beam-steering with time-modulated metasurfaces in the near-infrared spectral regime”
[…]
3. **Article Title:** ”Path Loss Modeling and Measurements for Reconfigurable Intelligent Surfaces in the Millimeter-Wave Frequency Band”
[…]
4. **Article Title:** ”Index Modulation for 6G Wireless Networks Using Time-Varying Metasurfaces”
[…]
From this analysis, we see that Qiang Cheng is an author in four different articles.
### Step 2: Count the Number of Articles Featuring Qiang Cheng
Now that we have identified the articles where Qiang Cheng is an author, we can count them. The list of articles includes:
1. ”Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design”
2. ”Broadband continuous beam-steering with time-modulated metasurfaces in the near-infrared spectral regime”
3. ”Path Loss Modeling and Measurements for Reconfigurable Intelligent Surfaces in the Millimeter-Wave Frequency Band”
4. ”Index Modulation for 6G Wireless Networks Using Time-Varying Metasurfaces”
[…]
### Final Answer
[…]
Question
Context 1M |
How many different articles has A. N. Pettitt contributed to? |
[…]

### v2

@@ Appendix B Experimental Settings
| Qwen3-4B-Instruct-2507 | 4B | ✓ | 262,144 | July 2025 |
| Qwen3-4B-Thinking-2507 | 4B | ✓ | 262,144 | July 2025 |
| Qwen3-30B-A3B-Instruct-2507 | 3Bx10E | ✓ | 262,144 | July 2025 |
| Qwen3-30B-A3B-Thinking-2507 | 3Bx10E | ✓ | 262,144 | July 2025 |
@@
| Full-text Articles | Database Tables | ||||||||
| Model | Context Size | F-64 | F-128 | F-512 | F-1024 | D-64 | D-128 | D-512 | D-1024 |
@@
| Qwen3-4B-Instruct-2507 | 256K | 7.2 | 7.3 | — | — | 47.2 | 46.0 | 38.9 | 31.1 |
| Qwen3-4B-Thinking-2507 | 256K | 50.2 | 42.3 | — | — | 92.7 | 85.8 | 79.4 | 63.6 |
| Qwen3-30B-A3B-Instruct-2507 | 256K | 15.0 | 16.4 | — | — | 55.6 | 53.8 | 44.7 | 42.0 |
| Qwen3-30B-A3B-Thinking-2507 | 256K | 62.5 | 55.1 | — | — | 94.4 | 88.3 | 80.8 | 74.5 |
@@
Appendix C Additional Results Using F1
Appendix D Details on Model Failure Analysis
To gain deeper insight into model behaviour, we analyzed the failure patterns exhibited in model-generated answers on SciTrek. We randomly selected 60 instances in the input length of 128K that were hard for all zero-shot and supervised models (10 instances per skill in Table 4). We find that models tend to output “NULL" particularly on samples that require the skill of filtering. As shown in Table 14, most open-weight models tend to randomly output “NULL" for any sample. This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context.151515Our instruction prompted models to generate “NULL” when they could not get the answer from the context. In addition, we find that models fail to follow the specified format for the answers, especially for sorting-related questions, as shown in Table 15. For example, models tend to generate author lists when sorted author numbers are requested and output lists when aggregates are requested. Interestingly, GPT-4.1 produces incorrectly formatted answers far more frequently than all other models. We further observe that some models tend to produce partial answers when lists are requested, especially for tasks that require aggregation. As shown in Table 16, our GRPO-based model tends to generate partial answers more than other models.
@@ Appendix D Details on Model Failure Analysis
We further manually analyzed 200 reasoning traces from reasoning models including Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates. We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis. For each SQL query template, we have different SQL queries and questions for different sizes of contexts, while the questions require the same reasoning capability. There are no reasoning traces when the input is 512K or 1M tokens for Qwen3-4B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B, as they do not support the input of 512K or 1M tokens.
| Model | A | S | F | F+A | F+S | RF |
@@
| Model | A | S | F | F+A | F+S | RF |
@@
| Model | A | S | F | F+A | F+S | RF |
@@
| Model | F+A | F+S | F | RF | Avg |
@@
Appendix E Prompts used for SciTrek Creation and Model Evaluation
We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language. Our prompt is presented in Figure 3. The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4. We employ two different prompts for (zero-shot) model evaluation. The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as context is shown in Figure 7. The prompt used by GRPO to generate reasoning races is given in Figure 6.